# Learning nuclear cross sections across the chart of nuclides with graph neural networks

### v2

@@ IV.2 GNN performance
The training of the GNN requires a partitioning of the TENDL dataset that is different from the one adopted in the training of the VAE and INR. Since the goal of the GNN, after training, is to make predictions of cross sections, possibly a few neutrons away from the training set, we divide the entire TENDL dataset into a training and test set based on a windowing function. The windowed test set is created using a rectangular window centered around a specific nucleus $(Z_{0},N_{0})$ in the chart of isotope and containing all nuclei with $Z_{0}-\Delta Z\leq Z_{0}\leq Z_{0}+\Delta Z$ and $N_{0}-\Delta N\leq N_{0}\leq N_{0}+\Delta N$. All results presented in this section were obtained for a test window centered on $(Z_{0},N_{0})=(42,60)$ with $\Delta Z=\Delta N=8$. This window contains a total of 289 test nuclei, 209 of which are included in the original TENDL dataset. Since that original dataset contains 2151 points on the 2D plane, this means that about 7.2% of the total data is allocated to the test set. Appendix VII.2 gives results obtained with different window center locations and different window sizes.
All of the GNN models are trained for $5000$ epochs with a learning rate of $1\times 10^{-3}$ on the training data. We use an Adam optimizer [42] with $\beta_{1}=0.9,\beta_{2}=0.999$. For each epoch, the entire training set is passed through the GNN and we do not create batches of the training graph since breaking the edge relationships between the nodes would yield inaccurate results. The prediction error of the decoded cross-sections is computed as
@@
where $\mathbf{\hat{x}}_{i}$ and $\mathbf{y}_{i}$ are the predicted node features and the ground truth in the original space, respectively, and $n$ is the actual number of nuclei in the set (either training or test). The quantity $(\mathbf{y}_{i}-\mathbf{\hat{x}}_{i})$ computes the point-wise difference and is a 256-dimension vector, which is then summed for each position $i$ in the test set. In the training phase $n=1942$ (the 2151 original nuclei of the TENDL dataset minus the 209 of the test window). During inference, the trained GNN is used to predict the node features in the test set and $n=209$. Since the model was only trained on the training data, there is no contamination bias from the test set on the GNN. The predicted cross-sections are then decoded using the chosen decoder $\mathcal{D}$, which is kept frozen to prevent any updates to the weights. We calculate the prediction error in the 256-dimension space using Eq. (8) and report it in Fig. 4 and Fig. 5.
In Fig. 4, the overall spread of the error is presented for all points in the dataset. We theorize that the traditional GCN architecture is too simple to estimate the structure of the TENDL grid. This is validated by the plots in Fig. 4 (bottom row), where the spread of the error in the GCN is quite large. On the other hand, the ARMA model shows excellent predictive performance in the test set for both the VAE and INR encoding. Interestingly, while the GatedConv model behaves as expected for the VAE encoding, we observe that for the case of the INR, there is a drop in performance. As we will discuss in Section V, this is a consequence of the fact that the INR latent space is more restricted than the one from the VAE.
Since the GNN seeks to estimate the latent manifold to learn the underlying relationships of the graph, a narrow distribution of the latent variables hinders the GNN from learning these relationships accurately. To better quantitatively measure the performance of each network, we list in Table 1 the average test PSNR using Eq. (7) over all points in the test set. We also provide comparisons with and without interpolation-based data augmentation, denoted as DA in the table.
@@
Table 1: Test PSNR values of the different GNN architectures with and without data augmentation (DA).
ARMA Gated GCN
VAE (DA) 37.80 $\mathbf{43.46}$ 27.75
VAE (no DA) 37.47 $\mathbf{43.44}$ 27.80
INR (DA) $\mathbf{42.23}$ 31.80 27.99
INR (no DA) $\mathbf{42.98}$ 31.63 28.10
Figure 5: Performance of the VAE and INR model for the GatedConv(left) and ARMAConv (right) GNN architecture in predicting nuclear cross-sections within the test set for the central nucleus, 102Mo and the ones at the lower left and upper right of the window, 86Se and 118Sn respectively. Plain curves without symbols correspond to the ground truth while dashed curves with open symbols correspond to GNN predictions. All cross sections are normalized independently in each channel across the entire dataset.
To better visualize the performance of each network in the prediction task, we show in Fig. 5 the normalized cross sections in three different nuclei within the test window: 86Se is at the “lower left” corner of the window; 102Mo is exactly at the center (hence the furthest away from the training set); 118Sn is at the “upper right” of the window. For each nucleus, we look at the performance of the GNN depending on whether the encoding was from a VAE or an INR. According to Table 1, the best GNN architecture with VAE encoding is GatedConv while it is ARMAConv for the INR encoding. For each encoding, the figure shows the results for the optimal GNN architecture. In the case of the VAE, the predictive power is rather good since most of the curves are hardly distinguishable, the exception being the $(n,2n)$ cross section in 86Se. The performance of ARMAConv with the INR encoding is visibly improved compared to that of the VAE, as reaffirmed by the values in Table 1. Note that the $(n,\gamma)$ channels is not identically zero but the normalized cross section is several order smaller than the other ones for these particular nuclei.
@@ V Discussion
The results presented in Section IV pose a few interesting questions, both from a data science and from a nuclear physics perspective, which we discuss in this section. In particular, we focus on a brief analysis of the structure of the latent space to analyze if specific patterns can be matched to nuclear properties. We also outline how our framework lends itself to propagating uncertainties of cross sections throughout the nuclear chart.
@@ V.1 Representation learning and structure of the latent space
First, we emphasize the importance of the training objective of Eq. (4) for successful GNN training. When training the GNN with the original cross-sections as targets (i.e., $\mathcal{L}_{\rm GNN}=\|\mathcal{D}(\chi(\mathcal{G}))-\mathbf{y}\|_{2}$), where $\mathcal{D}$ represents the VAE or INR decoder, and $\mathbf{y}$ is the original cross-sections with 256 dimensions, we observe that the GNN prediction performance is degraded with INR-based encoding compared to VAE. Unlike VAEs, which promote continuous representations through regularization (e.g., standard normal Gaussian prior), the INR’s latent space is more tightly constrained by its regularization term, particularly $\mathcal{L}_{\rm latent}$. This restriction may suppress features that would be discriminative for downstream prediction tasks. Additionally, the INR’s interconnected decoder architecture, where a hypernetwork generates the INR’s weights, could amplify the impact of perturbations in INR weights during GNN training, further hindering prediction performance.
@@
Figure 7 shows the two-dimensional representation of this t-SNE embedding for both the VAE and INR encodings as obtained with the scikit-learn implementation. Like Fig. 6, the color of a point is the difference $N-Z$ of the neutron number $N$ and proton number $Z$ in the corresponding nucleus. In practice, t-SNE algorithms require setting a few hyperparameters, most notably one called “perplexity” ${\tt Perp}$ which can be interpreted as “a smooth measure of the effective number of neighbors” in the high-dimensional space [43]. Although traditional implementations of t-SNE suggest values of the perplexity between 5 and 50, this is highly dependent on the dataset and its dimensionality. In our case, the dataset has 1936 points – training nuclei represent 90% of the whole TENDL dataset of 2151 nuclei – and is encoded in a latent space of dimension $k=32$. We found that in order to converge the Kullback-Leibler loss $\mathcal{L}_{PQ}$, the perplexity parameter had to be significantly larger: figure 7 was obtained for ${\tt Perp}=200$. Note that the spatial arrangement of points and the variability scale in each latent dimension are very dependent on the value of ${\tt Perp}$. In particular, small values of ${\tt Perp}$ (corresponding to large values of $\mathcal{L}_{PQ}$) lead to a lot of clustering and more compact latent spaces.
@@
To better analyze clustering structures, we used the scikit-learn implementation of the spectral clustering algorithm to identify potential clusters of points directly in the VAE or INR dataset. Out of all available clustering techniques, spectral clustering and the DBScan algorithms are the naive best choices to identify clusters based on existing documentation and literature. However, calibrating the maximum distance between any two points to be considered in the neighborhood of each other in DBscan requires a deeper understanding of the latent space than we currently have. In contrast, spectral clustering does not involve any prior knowledge of the latent space. We first determined a rough estimate of the optimal number $n_{\rm clus}$ of clusters based on the Silhouette method. In practice, we found $n_{\rm clus}=4$. Figure 8 shows the position in the nuclear chart of each nucleus from the INR encoding training dataset, color-coded by the cluster the nucleus belongs to. The most striking feature of this plot is the fact that nuclei are organized in diagonal bands loosely following the valley of stability rather than along isotopic, isotonic or isobaric sequences. In addition, some of the clusters seem to be related to neutron magic numbers (but, interestingly, not proton magic numbers), especially $N=82$ and $N=126$. Both observations are quite robust. In Appendix VII.3, we show that different clustering scenarios lead to similar qualitative results.
@@ V.2 Predictions of unknown cross sections
Predicting nuclear cross sections across the entire chart in a consistent manner and with well-defined uncertainties is particularly challenging due to the lack of experimental data on radioactive isotopes and the limited predictive power of existing models. In Section IV we have shown that both VAE and INR networks can successfully compress four-channel cross-sections with full energy index into compact, low-dimensional representations. This approach exhibits excellent encoding performance. This non-trivial result suggests that the “effective” space of nuclear cross sections is low-dimensional indeed.
The predictive ability of GNNs is dependent on the structure of the data, as observed in a supplementary experiment where we randomly sampled the TENDL graph to generate the training and test set (instead of the windowing function). In this setting the overall error across all models was higher by $\approx$ 30%. Additionally, as discussed previously, the structure of the latent space plays an important role in the optimization choice for the GNN. From Table 1 we see that in the case of the INR, while the ARMAConv model performs as expected, the GatedConvmodel performs worse compared to the VAE-encoding. It is important to note that the structure of the ARMAConv layer accounts for higher-order neighborhood information in the graph compared to the traditional polynomial layers used in GatedConv and GCNConv. Since ARMA filters can capture the global graph structure with a small number of learnable parameters ($\approx$ 250k), it can better estimate the “islands” in the INR hypernet latent space hence providing excellent predictive ability for the test set as seen in both Fig. 4 and Fig. 5. The GatedConv model does not have this advantage and hence we observe a drop in performance for the INR-encoding.
@@
V.3 Propagation of Uncertainties
To illustrate the power of our deep-learning model, we have performed the following “experiment”. We consider the $(n,2n)$ cross section in the 75As nucleus, which we note $\sigma_{i}(E)$, where $i$ simply indexes the position of 75As in the TENDL dataset. We generate 100 random realizations of the $\sigma_{i}(E)$, which we denote $\sigma^{(\alpha)}_{i}(E)$, by considering the covariance information listed in [29], which treats the entire evaluation as a large multivariate normal distribution. We use the VAE encoder to transform each $\sigma^{(\alpha)}_{i}(E)$ into a vector $\mathbf{x}_{i}^{(\alpha)}$ of the latent space. In the next step, we use these latent vectors $\mathbf{x}_{i}^{(\alpha)}$ to define new feature matrices. Keeping the notations of Section III.3, we write $\mathcal{G}^{(\alpha)}=(\mathcal{V}^{(\alpha)},\mathcal{E}^{(\alpha})$ the graph made of nodes $\mathcal{V}^{(\alpha)}$ and edges $\mathcal{E}^{(\alpha)}$. The set of all nodes $\mathcal{V}^{(\alpha)}$ depend on the feature matrix $\mathbf{X}^{(\alpha)}$ associated with the variation $\alpha$. Note that the only differences between any $\mathbf{X}^{(\alpha)}$ and $\mathbf{X}^{(\beta)}$ is the single latent vector corresponding to 75As: $\mathbf{x}_{i}^{(\alpha)}\neq\mathbf{x}_{i}^{(\beta)}$. However, the trained GNN is the set of parameters that define the function $\chi$ in Eq. (4). This means that the predictions from the trained GNN for each $\mathcal{G}^{(\alpha)}$ are all different from one another: the network is unchanged but its input (=the graph $\mathcal{G}^{(\alpha)}$) is.
Let us note $\Sigma_{i}(E)$ the cross section predicted by the GNN in nucleus $i$ based on the graph $\mathcal{G}$. Technically, $\Sigma_{i}(E)$ is obtained by running the decoder of the VAE or INR on the latent vector predicted by the GNN for that nucleus $i$. Since the GNN is not perfect, $\Sigma_{i}(E)\neq\sigma_{i}(E)$. Similarly, we note $\Sigma_{i}^{(\alpha)}(E)$ the cross section predicted by the GNN in the same nucleus $i$ but with the input graph $\mathcal{G}^{(\alpha)}$. We can now quantify the difference in GNN predictions from any arbitrary realization $\alpha$ by computing the deviation
$\varepsilon_{i}^{(\alpha)}=||\Sigma_{i}-\Sigma_{i}^{(\alpha)}||^{2}.$ (9)
The total mean deviation across the entire TENDL dataset is then just given by $\tfrac{1}{n}\sum_{i}^{n}\varepsilon_{i}^{(\alpha)}$.
Figure 9: Deviation (9) between the GNN prediction based on the initial set of TENDL cross sections and the one based on variation $\alpha=1$, for each nucleus in the TENDL dataset. The plot is normalized between $\epsilon_{\rm min}$ and $0.1\times\epsilon_{\rm max}$, where $\epsilon_{\rm min}$ ($\epsilon_{\rm max}$) is the minimum (maximum) error. The inset corresponds to the blue square centered at $Z=33$ and $N=42$.
In Figure 9, we plot the deviation (9) of the GNN for each nucleus in the TENDL dataset for the variation $\alpha=1$; similar results would be obtained for all other variations. In this experiment, the deviation should be thought of as a proxy to visualize the propagation of uncertainties. It shows how changes of one cross section in one nucleus propagate to neighboring nuclei. As visualized in the inset of the figure, changes are localized near the central nucleus 75As and decay rapidly as we go away from this nucleus.
Figure 10: Deviation (9) in 75As between prediction from the initial graph $\mathcal{G}$ and three linear combinations of the first random realization $\mathcal{G}^{(1)}$; see text for details. The difference is plotted in real units (barns).
Interestingly, changes do no propagate in a linear manner through the GNN. In Fig. 10, we plot the deviation (9) by generating three different graphs from the $\alpha=1$ variation cross section. Instead of replacing the latent vector $\mathbf{x}_{i}$ by the full variation $\mathbf{x}_{i}^{(\alpha)}$, we first introduce “damped” variations of the cross section characterized by the parameter $\lambda$, $\sigma_{i}\rightarrow\sigma_{i}^{(\alpha,\lambda)}=\sigma_{i}+\lambda(\sigma_{i}^{(\alpha)}-\sigma_{i})$. We then encode these $\sigma_{i}^{(\alpha,\lambda)}$ cross sections to define the corresponding feature matrix $\mathbf{X}^{(\alpha,\lambda)}$ before running the GNN and decoding its prediction. Figure 10 was obtained for three different values $\lambda=0.1,0.5,1.0$ of the damping factor. We see that the difference with the original prediction corresponding to the full variation ($\lambda=1.0$) is in fact smaller than for the other two cases.. The reason for this behavior is that each GNN learns an unique mapping function to predict the feature vector at a specific node based on the information from the neighboring nodes. Hence, the way that information is propagated depends on the nature of the mapping function, which is not always linear. However, our method still provides crucial insight into the behavior of neighboring nuclei surrounding the perturbation site.
Figure 11: Predictions of the $(n,2n)$ cross section in 74Ge ($\delta Z=-1$), 75As ($\delta Z=0$), and 76Se ($\delta Z=+1$) for the GNN with the initial graph $\mathcal{G}$ (plain lines, full symbols) and the full variation $\mathcal{G}^{(1)}$ (dashed lines, open symbols). Here, $\delta Z$ refers to the difference in proton number relative to 75As. Cross sections are plotted in real units (barns).
Finally, to visualize the concrete impact of these variations on actual cross sections, we plot in Fig. 11 the full $(n,2n)$ cross section as predicted by the GNN, either with the initial graph $\mathcal{G}$ – marked as ’(ini.)’ – or with the $\alpha=1$ variation graph $\mathcal{G}^{(1)}$ – marked as ’(var.)’. The plots are made for the central nucleus where the variation is generated, 75As, as well as for its two nearest neighbors along the proton number axis, 74Ge ($Z=32$ and $N=42$) and 76Se ($Z=34$ and $N=42$).
Running the GNN on such an updated set of cross sections – in our case, the $\mathcal{G}^{(\alpha)}$ for the variations $\alpha$ of the $(n,2n)$ cross section in 75As – takes less than a second on a laptop. It provides an estimate of how changes in one cross section impacts all other cross sections in all other nuclei. Crucially, this information is completely agnostic: whether the changes in that cross section come from new measurements, better theories, better evaluations or a combination of the above, does not matter. Our approach is completely data driven. As mentioned in Section II, it is simply impossible with the current state of nuclear theory and nuclear data evaluations to obtain the same information.

### v3

@@ IV.2 GNN performance
To evaluate the GNN performance, we divide the entire TENDL dataset into a training and testing set based on a windowing function. The windowed test set is created using a square window centered around a specific nucleus $(Z_{0},N_{0})$ in the chart of isotopes and containing all nuclei with $Z_{0}-\Delta Z\leq Z_{0}\leq Z_{0}+\Delta Z$ and $N_{0}-\Delta N\leq N_{0}\leq N_{0}+\Delta N$. All results presented in this section were obtained using a test window centered at $(Z_{0},N_{0})=(42,60)$ with $\Delta Z=\Delta N=4$, resulting in a $9\times 9$ test region on the nuclear chart. This window contains a total of 81 test nuclei, 67 of which are included in the original TENDL dataset. Since that original dataset contains 2171 points on the 2D plane, this means that about 3.0% of the total data is allocated to the test set. In Appendix VII.2, we also present the GNN performance across varying window sizes and different center locations.
Figure 5: Performance of the VAE (left) and INR (right) model for the GraphSAGE GNN architecture in predicting nuclear cross sections within the test set for the central nucleus, 102Mo and the ones near the lower left and upper right of the window, 95Y and 108Ru respectively. Plain curves without symbols correspond to the ground truth while dashed curves with open symbols correspond to GNN predictions. All cross sections are normalized independently in each channel across the entire dataset. The small grid in the upper right corner of the INR figures shows the location of each nucleus within the test set.
To train the GNNs, we use the AdamW optimizer [42] with a base learning rate of $lr=0.001$. To promote faster convergence and stabilize training dynamics, we adopt a OneCycle learning rate scheduler [43], which linearly increases rate up to $10\times lr$ before annealing it back down. Specifically, we set the maximum learning rate to $0.01$, with a total of $100,000$ training epochs. This setup was applied consistently across both VAE-based and INR-based experiments. For each epoch, the entire training set is passed through the GNN and we do not create batches of the training graph since breaking the edge relationships between the nodes would yield inaccurate results. The prediction error of the decoded cross sections is computed as:
@@
where $\mathbf{\hat{x}}_{i}$ and $\mathbf{y}_{i}$ are the predicted node features and the ground truth in the original space, respectively, and $n$ is the actual number of nuclei in the set (either training or test). The quantity $(\mathbf{y}_{i}-\mathbf{\hat{x}}_{i})$ computes the point-wise difference, which is then summed for each position $i$ in the test set. We calculate the prediction error in the 256-dimension space using Eq. (9), and present the corresponding error distribution as a heat map in Fig. 4.
In Fig. 4, we observe notable differences in the error distribution between VAE and INR embeddings across the nuclear chart. In the VAE case (left column), prediction errors are mostly concentrated in the test region, while the rest of the chart (training region) shows low errors. Conversely, INR-based models (right column) display more widespread errors, including in the training region, especially in the upper part of the chart. This aligns with INR’s relatively compact latent space as well as lower diversity across samples, which leads to a smaller training-testing gap. We further examine and visualize the characteristics of these latent spaces in Section V.1. To provide a more quantitative comparison, Table 1 reports the average test PSNR of Eq. (8) over all test samples for each GNN architecture using both VAE and INR embeddings.
@@
Table 1: Test PSNR values of the different GNN architectures with VAE and INR embeddings.
GNN Arch. ARMAConv GatedConv GraphSAGE
INR embedding 35.31 35.48 35.14
VAE embedding 35.87 36.40 $\mathbf{36.65}$
We also visualize the prediction performance of the GraphSAGE network using both VAE-encoding (left column) and INR-encoding (right column) in Fig. 5. The figure shows the normalized cross sections for three nuclei located within in the test window: 95Y near the lower-left corner, 102Mo at the center (hence the furthest away from the training set), and 108Ru near the upper-right corner of the window. With the VAE-encoding, the predicted curves closely match the ground truth, with only a slight deviation in the $(n,2n)$ channel. The predictions from the INR-encoding exhibit similar trends and capture the overall structure of the cross sections, despite more noticeable differences in $(n,2n)$ channel, especially at the center region. Numerical experiments show that the model’s epistemic error stemming from the training of both VAE/INR and GNN is small and have no effect on these conclusions. Note that the $(n,\gamma)$ channels is not identically zero but the normalized cross section is several order smaller than the other ones for these particular nuclei.
@@ V Discussion
The results presented in Section IV pose a few interesting questions, both from a data science and from a nuclear physics perspective, which we discuss in this section. In particular, we focus on a brief analysis of the structure of the latent space to analyze whether specific patterns can be matched to nuclear properties.
@@ V.1 Representation learning and structure of the latent space
When training GNNs with different latent representations, the choice of the loss function (6) or (7) significantly impacts performance. Our experiments reveal a clear contrast: GNNs trained on VAE-based embeddings achieve better performance when optimizing in the original cross section space, whereas GNNs trained on INR-based embeddings perform better when optimized directly in the latent space. This discrepancy can be attributed to two key factors:
Decoder sensitivity in INR models
The INR decoder is parameterized by a hypernetwork which maps latent vectors not directly to outputs but to the weights of an implicit neural function. This architectural design introduces an additional sensitivity when it is used as downstream prediction task. Small perturbations in the predicted weights of the neural representation can drastically affect the predicted cross sections. This makes finetuning the decoder in an end-to-end fashion less effective compared to the one with VAE-based embedding.
Embedding space expressivity
As shown in Fig. 6, VAE latent spaces tend to have higher variance and broader spread across samples, which provide the downstream GNN with richer signals to learn from, whereas, INR latent spaces are more compact and concentrated in narrow regions of the latent space. This could limit the discriminative power of the GNN. Consequently, optimizing directly in the latent space for INR-based embedding circumvents the difficulty of optimization through the sensitive decoder.
@@
Figure 7 shows the two-dimensional representation of this t-SNE embedding for both the VAE and INR encodings as obtained with the scikit-learn implementation. Like Fig. 6, the color of a point is the difference $N-Z$ of the neutron number $N$ and proton number $Z$ in the corresponding nucleus. In practice, t-SNE algorithms require setting a few hyperparameters, most notably one called “perplexity” ${\tt Perp}$ which can be interpreted as “a smooth measure of the effective number of neighbors” in the high-dimensional space [44]. Although traditional applications of t-SNE suggest values of the perplexity between 5 and 50, this is highly dependent on the dataset and its dimensionality. In our case, the dataset has 2090 points—training nuclei represent 97% of the whole TENDL dataset of 2171 nuclei—and is encoded in a latent space of dimension $k=32$. We found that in order to converge the Kullback-Leibler loss $\mathcal{L}_{PQ}$, the perplexity parameter had to be significantly larger: Fig. 7 was obtained for ${\tt Perp}=200$. Note that the spatial arrangement of points and the variability scale in each latent dimension are very dependent on the value of ${\tt Perp}$. In particular, small values of ${\tt Perp}$ (corresponding to large values of $\mathcal{L}_{PQ}$) lead to excessive visible clustering while failing to converge the projection.
@@
To better analyze clustering structures, we used the scikit-learn implementation of the spectral clustering algorithm to identify potential clusters of points directly in the VAE or INR dataset. Out of all available clustering techniques, spectral clustering and the DBScan algorithms are the naive best choices to identify clusters based on existing documentation and literature. However, calibrating the maximum distance between any two points to be considered in the neighborhood of each other in DBscan requires a deeper understanding of the latent space than we currently have. In contrast, spectral clustering does not involve any prior knowledge of the latent space. We first determined a rough estimate of the optimal number $n_{\rm clus}$ of clusters based on the Silhouette method. In practice, we found $n_{\rm clus}=4$. Figure 8 shows the position in the nuclear chart of each nucleus from the INR encoding training dataset, color-coded by the cluster the nucleus belongs to. The most striking feature of this plot is the fact that nuclei are organized in diagonal bands loosely following the valley of stability rather than along isotopic, isotonic or isobaric sequences. In addition, some of the clusters seem to be related to neutron magic numbers (but, interestingly, not proton magic numbers), especially $N=82$ and $N=126$. Both observations are quite robust. In Appendix VII.4, we show that different clustering scenarios lead to similar qualitative results.
@@ V.2 Predictions of unknown cross sections
Predicting nuclear cross sections across the entire chart in a consistent manner and with well-defined uncertainties is particularly challenging due to the lack of experimental data on radioactive isotopes and the limited predictive power of existing models. In Section IV we have shown that both VAE and INR networks can successfully compress four-channel cross sections with full energy index into compact, low-dimensional representations. This approach exhibits excellent encoding performance. This non-trivial result suggests that the “effective” space of nuclear cross sections is low-dimensional indeed.
The predictive performance of GNNs in our imputation task is influenced by several interconnected factors. One of the most critical elements is the test window size, which directly impacts the number of available training nodes in the transductive setting where the model is trained and evaluated on a fixed global graph. As the size of the test window increases, fewer nodes remains for training, thereby limiting the opportunity for the model to learn from surrounding neighborhoods. This degradation in performance with larger test windows is confirmed in Fig. 11. Furthermore, nuclei occupy a diagonal band in the $(Z,N)$ chart rather than a uniform 2D grid. This results in regions of sparsity and disconnection that challenge the GNN’s ability to propagate information effectively. As a result, the location of the test window also significantly affects performance, for example, regions with denser nuclear neighborhood exploit more effective message passing, whereas sparse areas hinder it as verified in Fig. 10. In terms of training strategy, our results show that GNN performance depends not only on the structure of the latent space learned through VAEs or INRs but also on how the GNN is optimized. Specifically, the GNN with VAE embeddings benefit from end-to-end training in the original cross section space, while INR embeddings perform better when optimization is restricted to the latent space as reported in Table 2. Our analysis from mulitple perspectives, including test PSNR, error heatmaps, and the structure of the learned latent space, highlights the critical importance of the embedding in determining the effectiveness of downstream GNN predictions. These findings suggest that developing latent representations that are well aligned with the needs of graph-based prediction tasks is a promising direction for future research.